OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (anonymous):

C =5/9(F-32) hence C*9/5=F-32 hence F=C*9/5 +32 Then, F=2*9/5 +32 =18/5 +32 = 3.6+32=35.6 Fahrenheit